# From Block Diagrams to Bloch Spheres: Graphical Quantum Circuit Simulation in LabVIEW

### v1

@@ Abstract
As quantum computing transitions from theoretical physics to engineering applications, there is a growing need for accessible simulation tools that bridge the gap between abstract linear algebra and practical implementation. While text-based frameworks (like Qiskit or Cirq) are standard, they often present a steep learning curve for students and engineers accustomed to graphical system design. This paper introduces QuVI (Quantum Virtual Instrument), an open-source quantum circuit toolkit developed natively within the NI LabVIEW environment. Moving beyond initial proof-of-concept models, QuVI establishes a robust framework that leverages LabVIEW’s “dataflow” paradigm—where wires represent data and nodes represent operations—to provide an intuitive, visual analog to standard quantum circuit notation while enabling the seamless integration of classical control structures like loops and conditionals. The toolkit’s capabilities are demonstrated through the construction and visualization of fundamental quantum algorithms, verifying results against theoretical predictions. By translating “Block Diagrams” directly into quantum state evolutions (“Bloch Spheres”), QuVI offers educators and researchers a powerful platform for prototyping quantum logic without leaving the graphical engineering workspace.
@@ I Introduction
LabVIEW (Laboratory Virtual Instrument Engineering Workbench) offers a unique alternative that bridges this gap. Its graphical programming language (G) is inherently parallel and isomorphic to quantum circuit diagrams: wires carry state information, and nodes perform operations. Unlike web-based simulators, LabVIEW is a full-featured programming environment. This allows QuVI users to wrap quantum circuits inside standard structures—such as For Loops, While Loops, and Case Structures—enabling the seamless construction of complex algorithms where classical logic controls quantum execution flow.
This work presents QuVI, a versatile quantum circuit toolkit for LabVIEW which is publicly available via the VI Package Manager. [^1] https://www.vipm.io/package/murtaza_vefadar_lib_quvi___quantum_circuit_toolkit/ A custom state-vector update engine is implemented to exploit CPU parallelism for both single-qubit controlled gates and multi-qubit operations, and the system is validated against standard quantum benchmarks.
@@ II.2 Circuit Composition
    Initialization: The circuit begins with ”QuVI Wire” subVIs, each initializing a qubit in the $|0\rangle$ state. These are stacked vertically to define the register size.
@@ II.3 Synchronization with Notifiers
Figure 1: QuVI architecture demonstrating execution flow. Top wires: Independent operations execute in parallel via native dataflow. Bottom wires: Entangling gates (e.g., CNOT) enforce synchronization using Notifiers to handle state dependencies between control and target qubits.
When a controlled gate is executed, the target gate first performs the necessary calculations on the global state vector. Once the update is complete, the Watch List bits are modified according to the control bitmask (inclusion and value masks). A LabVIEW Notifier is then employed to signal the waiting gates situated immediately after the control and anti-control nodes. Upon receiving this notification, these gates access the updated Watch List. Execution for a specific wire proceeds only if the corresponding bit in the array is observed to be “off,” allowing the gate to safely begin its own operations on the state vector.
@@ IV.2 Grover’s Search Algorithm
The core sequence of the algorithm—the Oracle ($O$) followed by the Diffusion operator (DD)—is placed inside a LabVIEW For Loop structure (see Figure 4). This allows the user to automate the amplitude amplification process, running the sequence $R\approx\frac{\pi}{4}\sqrt{N}$ times (here $R=3$) without manually duplicating the block diagram components. Additionally, a Case Structure is used to dynamically select the Phase Oracle based on a user input (here set to tagged state $|0101\rangle$), allowing a single VI to search for any target state.
The system was initialized in the equal superposition state $|s\rangle=H^{\otimes 4}|0\rangle$. After 3 iterations determined by the For Loop, the simulator correctly amplified the probability of measuring the dynamically selected target state ($|0101\rangle$) to $96\%$, consistent with theoretical predictions.
Figure 4: 4-Qubit Grover Search in QuVI. (a) LabVIEW Block Diagram utilizing a For Loop for algorithm iterations and a Case Structure for dynamic Oracle selection. The tagged state is set to $|0101\rangle$. (b) Simulation result showing the probability distribution of the final state, with a dominant peak ($96\%$) at the index 0101 (decimal 5).
@@ V Conclusion
Future enhancements will focus on extending the simulator beyond pure state vectors to support mixed states via the density matrix formalism. This development will enable the simulation of real, noisy quantum systems through the incorporation of decoherence channels. Furthermore, diagnostic tools for quantifying entanglement—such as von Neumann entropy and concurrence [6]—and computing reduced density matrices via partial traces [8] will be implemented.

### v2

@@ Abstract
As quantum computing transitions from theoretical physics to engineering applications, there is a growing need for accessible simulation tools that bridge the gap between abstract linear algebra and practical implementation. While text-based frameworks (like Qiskit or Cirq) are standard, they often present a steep learning curve for students and engineers accustomed to graphical system design. This paper introduces QuVI (Quantum Virtual Instrument), an open-source quantum circuit toolkit developed natively within the NI LabVIEW environment. Moving beyond initial proof-of-concept models, QuVI establishes a robust framework that leverages LabVIEW’s “dataflow” paradigm, in which wires represent data and nodes represent operations, to provide an intuitive, visual analog to standard quantum circuit notation while enabling the seamless integration of classical control structures like loops and conditionals. The toolkit’s capabilities are demonstrated by constructing and visualizing fundamental quantum algorithms and verifying results against theoretical predictions. By translating “Block Diagrams” directly into quantum state evolutions (“Bloch Spheres”), QuVI offers researchers and educators a powerful platform for prototyping quantum logic without leaving the graphical engineering workspace.
@@ I Introduction
LabVIEW (Laboratory Virtual Instrument Engineering Workbench) offers a unique alternative that bridges this gap. Its graphical programming language (G) is inherently parallel and isomorphic to quantum circuit diagrams: wires carry state information, and nodes perform operations. Unlike web-based simulators, LabVIEW is a full-featured programming environment.
This work presents QuVI, a versatile quantum circuit toolkit for LabVIEW, which allows its users to wrap quantum circuits inside standard structures (such as For Loops, While Loops, and Case Structures), enabling the seamless construction of complex algorithms where classical logic controls quantum execution flow. A custom state-vector update engine is implemented to exploit CPU parallelism for both single-qubit controlled gates and multi-qubit operations, and the system is validated against standard quantum benchmarks. QuVI is publicly available via the VI Package Manager  [^1] https://www.vipm.io/package/murtaza_vefadar_lib_quvi___quantum_circuit_toolkit/.
@@ II.2 Circuit Composition
    Initialization: The circuit begins with ”QuVI Wire” subVIs, each initializing a qubit in the $|0\rangle$ state. These are stacked vertically to define the register size, following the convention that the topmost wire corresponds to $q_{0}$.
@@ II.3 Synchronization with Notifiers
Figure 1: QuVI architecture demonstrating execution flow. Top wire: independent operations execute in parallel via native dataflow. Bottom wires: entangling gates (e.g., CNOT) enforce synchronization using Notifiers to handle state dependencies between control and target qubits.
When a controlled gate is executed, the target gate first performs the necessary calculations on the global state vector. Once the update is complete, the Watch List bits are modified according to the control bitmask (”inclusion mask” in the next section). A LabVIEW Notifier is then employed to signal the waiting gates situated immediately after the control and anti-control nodes. Upon receiving this notification, these gates access the updated Watch List. Execution for a specific wire proceeds only if the corresponding bit in the array is observed to be “off,” allowing the gate to safely begin its own operations on the state vector.
@@ IV.2 Grover’s Search Algorithm
The core sequence of the algorithm, the Oracle ($O$) followed by the Diffusion operator ($D$), is placed inside a LabVIEW For Loop structure (see Figure 4). This allows the user to automate the amplitude amplification process, running the sequence $R\approx\frac{\pi}{4}\sqrt{N}$ times (here $R=3$) without manually duplicating the block diagram components. Additionally, a Case Structure is used to dynamically select the Phase Oracle based on a user input (here set to tagged state $|0110\rangle$), allowing a single VI to search for any target state.
The system was initialized in the equal superposition state $|s\rangle=H^{\otimes 4}|0\rangle$. After 3 iterations determined by the For Loop, the simulator correctly amplified the probability of measuring the dynamically selected target state ($|0110\rangle$) to $96\%$, consistent with theoretical predictions.
Figure 4: 4-Qubit Grover Search in QuVI. (a) LabVIEW Block Diagram utilizing a For Loop for algorithm iterations and a Case Structure for dynamic Oracle selection. The tagged state is set to $|0110\rangle$. (b) Simulation result showing the probability distribution of the final state, with a dominant peak ($96\%$) at the index 0110.
@@ V Conclusion
Future enhancements will focus on extending the simulator beyond pure state vectors to support mixed states via the density matrix formalism. This development will enable the simulation of real, noisy quantum systems through the incorporation of decoherence channels. Furthermore, diagnostic tools for quantifying entanglement, such as von Neumann entropy and concurrence [6], and computing reduced density matrices via partial traces [8] will be implemented.